### v1

# Analysis of Josephson Junction Barrier Variation - a Combined STEM, Breakdown and Monte-Carlo Approach

Abstract
Josephson junctions manufactured to tight tolerances are necessary components for superconducting quantum computing. Developing precise manufacturing techniques for Josephson junctions requires an understanding of their make-up and robust feedback metrics against which to optimise. Here we consider complementary techniques assessing what conclusions they allow us to draw about the barriers in junctions. Monte-Carlo simulations of barriers show that standard deviations of 15-20% of the total barrier thickness are compatible with our experimental data. Electrical breakdown allows us to probe the weakest points in barriers. Narrowing the distribution of this breakdown provides a promising feedback mechanism for barrier optimisation. Grouping junctions by breakdown voltage allows us to identify sub-ensembles of junctions with different median resistance, highlighting the importance of weak spots in junction parameters such as resistance. Transmission electron microscopy can be used to find average barrier thickness, although we highlight challenges forming robust conclusions on the distribution of thicknesses in a barrier from these experiments.
preprint: APS/123-QED
I Introduction
Josephson junctions (JJs) are ubiquitous to superconducting quantum computers [1] and other superconducting circuits such as parametric amplifiers [2] and therefore have significant technological importance. The most widely used type of JJ in superconducting technologies is based on an insulating layer of AlOx which separates two superconducting electrodes. The AlOx layer is formed by oxidising the surface of a layer of metallic aluminium before covering the AlOx layer with a second superconducting electrode, also aluminium. This creates a barrier which is (i) amorphous (ii) non-stoichiometric, and (iii) with thickness that varies across the junction [3, 4, 5, 6]. Developing manufacturing techniques for JJs to be made with fine tolerances requires a detailed understanding of the junction, as well as robust feedback metrics against which manufacturing techniques can be optimised.
The figure of merit of an ideal JJ is its critical current, $I_{\rm C}$, a property that can be hard to measure at numbers that are statistically significant. A proxy for this value is the room temperature resistance [7, 8] and resistance tolerances are typically used to characterise manufacturing processes. These processes are in turn optimised considering feedback from resistance measurements and techniques such as scanning electron microscopy [9, 10]. Consequently many problems with repeatable JJ manufacture have been solved and control of the area of JJs is now much improved [7, 11]. As a result the current state of the art in as-fabricated JJ resistance spread is $\sim$2-3[12, 13, 14]. Once lateral dimensions of junctions are perfectly controlled the remaining variation between junctions must be attributed to differences between their barriers although it is not clear what level of resistance variation this causes in typical manufacturing processes.
[…]
Variation in dielectric thin films, like the barrier, is a well studied problem with relevance to devices including field effect transistors (FETs) [19], metal-insulator-semiconductor (MIS) or metal-insulator-metal (MIM) structures [20, 21] and magnetic tunnel junctions (MTJs) [22]. The barriers in JJs are typically $\sim 1-2$ nm [23] so even a single O2- ion represents a substantial fraction of the average barrier thickness with ionic radius of $\sim$0.14 nm [24]. Generating a full 3D map of barriers at this resolution isn’t yet possible although various techniques have been used to improve the understanding of barriers. (Scanning) Transmission Electron Microscopy ((S)TEM) and associated techniques give chemical and structural information at sub-nanometre resolution [4, 25]. Atom Probe Tomography (APT) results in 3D reconstructions of the barrier although to date, results on AlOx barriers do not show the requisite resolution [26, 25]. Scanning Tunneling Microscopy (STM) maps the as-grown oxide prior to being capped by a final metal layer [22] and oxygen migration occurs after this final capping [27]. Simulations of the growth process by molecular dynamics give insight into atomic structures in typical barriers [27]. Studies of breakdown voltages of barriers are used in device communities, such as FETs, to probe device-to-device variation in barrier weak points [28] although have not seen substantial use in JJs.
In this work we consider electrical measurements and STEM imaging of barriers, discussing what conclusions can be robustly drawn from these types of characterisation. We extend the typical range of electrical measurements to sample non-linear parts of the current-voltage (IV) curves and junction breakdown. We complement this by Monte-Carlo simulations of IVs and breakdown resulting from non-uniform barrier thickness. We are able to recreate the extended measurements with these Monte-Carlo simulations considering only statistically distributed thicknesses within single junctions and find that skewed thickness distributions fit our experiments best. We then consider how thickness information is derived from STEM imaging, a technique that has previously been used to measure thickness distributions in barriers. We consider how choices in image processing and non-idealities in the barrier electrodes may affect conclusions drawn from this type of analysis.
II Electrical measurements
Fig. 1 shows results from fitting a set of 598 IV curves. These are collected from junctions manufactured with the same target dimensions. An atomic force microscope (AFM) measurement of one of these junctions is shown in Fig. 1 (a). For each junction the voltage is swept from 0 V to a value up to 1.6 V as shown in Fig. 1 (b). At low voltages the IVs of the JJs are linear and we fit this dependence at voltages below 0.02 V to extract the junction resistance. At voltages below 1.6 V, each junction fails at its breakdown voltage and becomes an ohmic channel with much lower resistance (typically hundreds of Ohms).
We also fit the IVs for all voltages below the breakdown voltage using the Simmons model, which gives the current-voltage properties for a thin rectangular tunnel junction as a function of the potential barrier height and its thickness [29];
$I=\frac{Ae^{2}}{2\pi ht^{2}}\Biggl{[}\left(\phi-\frac{V}{2}\right)\exp\left(-K\sqrt{\phi-\frac{V}{2}}\right)-\\ \left(\phi+\frac{V}{2}\right)\exp\left(-K\sqrt{\phi+\frac{V}{2}}\right)\Biggr{]}$ (1)
where $A$ is the junction area, $t$ the junction thickness, $\phi$ the barrier height and $K=4\pi t\sqrt{2m_{e}e}/h$. Fitting this model gives a single-valued thickness which, as discussed in the introduction and shown in the literature, for example by STM [22], is a simplification of the true barrier in these junctions. The simple analytical form makes it computationally inexpensive to compute, which is important when used in Monte-Carlo simulations shown later.
[…]
We perform Monte-Carlo simulations where we allow thickness to vary across a junction trying to understand if we can recreate experimental measurements considering only thickness variations within junctions. To build these simulations we define a junction with an area corresponding to the overlap between the two electrodes in a JJ (shown in Fig. 1 a and the supplemental materials). This area is divided into equally sized pixels which act as parallel conduction channels. The thickness of the barrier at each pixel is drawn from a statistical distribution and then Eq. 1 is computed as the IV for that pixel. Combining the IVs of each pixel as parallel conduction paths, simulates the IV of the whole junction. We consider both normally distributed and log-normally distributed thicknesses, where a log-normal distribution is an example of a skewed thickness distribution.
Figure 1: (a) An AFM measurement of a JJ from the ensemble of JJs (b) An IV of a typical junction showing the extraction of the breakdown voltage, resistance from a linear fit to the low voltage regime and a fit to the full Simmons model. (c,d) Results from analysing an ensemble of 598 junctions fabricated on a 3” wafer. (c) A histogram showing the outputs from fitting Eq. 1 to measured IVs with thickness, barrier height and nominal area shown on the same x axis. (d) A histogram of the resistance deviation from the median resistance of the junction ensemble with a median resistance value of 7122 $\Omega$.
Figure 2: Results from Monte-Carlo simulations of barriers with normally (a - e) and log-normally (f - j) distributed thicknesses. Heatmaps of (a, f) resistance, (b, g) resistance spread (c, h) barrier height ($\phi$), (d, i) fitted thickness and (e, j) fitted area. Resistance and its spread are found by fitting a straight line to the low voltage region and other parameters are from found from fitting Eq. 1 to the full voltage range. Any regions within the experimental variation (or 10% instead for resistance) are coloured grey as shown in the colour bars. Overlaid are ‘bulls-eye’ concentric circle markers, which indicate a set of parameters which satisfies all the constraints set by the experimental measurements.
We characterise each simulated IV curve by refitting it to Eq 1 which allows us to compare simulated IVs to experimental IVs. Fitting the linear regime of the IV gives the resistance of the junction. Repeating this protocol for multiple junctions (20 in this work) gives the junction-to-junction variation for a given thickness distribution. The pixel size is an important parameter in these simulations with smaller pixels giving less junction-to-junction variation. Pixels less than $\sim$0.2 nm become unphysical as they drop below ionic radii of the barrier giving a lower bound on pixel size. We perform these simulations with a pixel size of 1 nm and a barrier height set to the median experimental value shown in Fig. 1 (b) $\phi=1.22$ eV. The pixel size is chosen to correspond approximately to the thickness of barriers assuming that the barrier thickness doesn’t change on a length-scale much less than the thickness and also as a compromise for computational speed. The value of barrier height is important for inferring accurate thicknesses from any subsequent modeling. Literature values vary with changes in barrier height being ascribed to the termination of the oxide barrier (Al/O) and the crystal structure of the underlying metal [30, 31]. Experimental studies also using the Simmons model identify barrier heights ranging from $\sim$0.9 eV to $\sim$1 eV [32]. Density functional theory (DFT) simulations of amorphous Al/AlOx/Al barriers show dependence on roughness of the underlying metal and find effective barrier heights ranging from 0.2 eV to 1 eV [33].
We show the results of these Monte-Carlo simulations in Fig. 2. The top and bottom rows show results for normally distributed and log-normally distributed barrier thicknesses respectively. In each heatmap we colour any cells compatible with our experimental results in grey. For both distributions there is a region where the experimental results shown in Fig. 1 are matched well by simulations indicated by ‘bulls-eye’ symbols in Fig. 2. Using a multi-valued thickness within a junction we are able to match experimental IVs whilst maintaining the known lateral dimensions of the junctions.
Figure 3: Comparison of (a) experimentally measured breakdown voltages with simulated breakdown voltages for (b) log-normally and (c) normally distributed barriers created with distributions indicated by stars in Fig. 2. (a) Shows a double-Gaussian fit indicating a bimodal breakdown distribution. Simulated distributions are found by taking the thinnest points in randomly generated barriers with a mesh size of 0.2 nm and rescaling this by a dielectric strength ($E_{\rm ds}$) to keep the mean breakdown voltage equal to the experimental values. (d) The resistance of junctions distinguishing junctions with breakdown above/below the midpoint of the bimodal distribution 1.3 V, shown by the black dashed line in (a). (e) A histogram of the full thickness distribution of example junctions generated with the parameters from (b, c). (f) Cumulative conductance of junctions from 10 different junctions generated with parameters from (e) as a fraction of the total junction area.
These simulations also show that inferring the barrier height by extracting the value from fits to the Simmons model is not straightforward. In Fig. 2 we use a single value of barrier height $\phi=1.22$ eV. However, the value of $\phi$ extracted from refitting these simulations shown in Fig. 2 (c,h) deviates from this value, depending on the underlying thickness distribution. For instance, moving along the x axis for fixed value of $\langle t\rangle$ in both Fig. 2 (c,h) the barriers with greater thickness variation appear to have a larger barrier heights. This may explain observations such as those seen in Ref. 32 where they conclude that oxidation time has changed the barrier height. Instead of changing barrier height, it may be that the oxidation time is changing thickness distributions (either average or variation), which appear as potential barrier heights when interpreted via fitting to the Simmons model. Given the variation in literature values and imperfect inference from electrical measurements, direct measurements of the barrier height would be advantageous, for instance by x-ray photoelectron spectroscopy techniques such as in Ref. [34].
To understand the effect of changing barrier height, we perform these simulations for $\phi$ = 1 eV and 0.8 eV and show these in the supplemental information and match the results of fitting the Simmons model to data (as indicated by the ‘bulls-eye’ in Fig. 2). With lower barrier heights, we infer a thicker average barrier and also a larger thickness standard deviation. When using the barrier height of 0.8 eV we find average thicknesses which matches our best inference from STEM imaging as discussed below. This places a lower bound on barrier heights compatible with our results. The use of a rectangular potential barrier will give a systematic offset implying a thinner barrier relative to what would be inferred using more realistic rounded potential barriers. A detailed understanding of potential barrier shapes or atomistic computations of tunneling currents would improve this model and likely result in better quantitative agreement between simulated potential barrier thicknesses and the thicknesses of alumina barriers. We could also extend this model to also draw $\phi$ from a distribution of values indicating compositional fluctuations at the expense of extra computational time. An understanding of realistic levels of potential barrier fluctuations would be necessary for this to give deeper insight to our results. The use of a single-valued potential-barrier height here is equivalent to an assumption that fluctuations in thickness contribute more to IVs than fluctuations in barrier height.
We can complement the information shown from the fitted IV curves with measurements of the breakdown voltage, the point indicated by a star in Fig. 1 (a). We present a histogram showing the analysis of junction breakdown voltages of the same group of junctions in Fig. 3 (a). We phenomenologically fit a double-Gaussian to the breakdown voltage showing a bimodal distribution with a midpoint of 1.3 V. Breakdown occurs at the weakest point in the barrier at a voltage given by $V_{\rm BD}=tE_{\rm ds}$ where t is the barrier thickness and $E_{\rm ds}$ its dielectric strength. Assuming a uniform dielectric strength across the barrier, the weakest point is determined by the thinnest point in the junction. Dielectric strength could also fluctuate across the barrier, for instance decreasing at defects in the barrier, where what constitutes a defect in an amorphous oxide is a complicated issue in itself [35]. If all junctions have the same thickness at their thinnest point, then the spread in Fig. 3 (a) would be a direct measure of the spread in dielectric strength, allowing us to place a bound on the standard deviation of this material property in these barriers, $\sigma_{E_{\rm DS}}<4$. We consider the case where dielectric strength is constant across all barriers and within each barrier, i.e. omitting any contributions from local structural changes. In this limit the breakdown voltage is determined by the thinnest point of each barrier and the histogram in Fig. 3 (a) is a rescaled measurement of these thinnest points from 598 junctions.
Extending the Monte-Carlo simulations shown in Fig. 2 we consider whether the same model of varying barrier thickness agrees with our experimental measurements of breakdown voltage. Using thickness distributions identified in Fig. 2 as giving the best match between simulated and experimental IVs we compute a distribution of breakdown voltages. We generate thicknesses from the distributions indicated by ‘bulls-eye’ symbols in Fig. 2 for a $240\times 240$ nm junction with a 0.2 nm pixel size 1110.2 nm approximates the ionic radii in the AlOx barrier. We record the minimum thickness of the barrier in the junction and then repeat this to build a distribution of the thinnest points in the barrier. In order to translate the distribution of the thinnest points to breakdown voltages we rescale the distribution of thin-points by a constant dielectric strength of 3.31 GV/m and 5.26 GV/m for log-normal and normal barrier distributions respectively. Previous measurements of $E_{ds}$ in amorphous alumina vary from 0.4 to 1.0 GV/m , similar in order of magnitude [37]. We compute the dielectric strength by enforcing the mean of the rescaled thicknesses to be equal to the mean of the measured breakdown voltages. Simulated breakdown-distributions for barriers with log-normally and normally distributed thickness are shown in Fig. 3 (b) and (c) respectively. Neither statistical distribution recreates the bimodal distribution seen experimentally. However, the rescaled values from log-normally distributed barriers (Fig. 3 b) show reasonably good agreement with the measured breakdown voltages with a similar standard deviation and min/max ratio. In contrast the rescaled values taken from barriers generated with a normally distributed thickness (Fig. 3 c) do not agree with the experimental measurements. We show that multiple experimental measurements can be largely recreated considering only a skewed thickness distribution in our barriers. The parameters used to simulate breakdown are drawn from simulations of IV curves which have no awareness of breakdown physics and only one value is used to rescale the distribution. We note we do not recreate a bimodal distribution indicating that, while better than a normal distribution, the log-normal thickness distribution is not a perfect description of the barrier.
In Fig. 3 (d) we present histograms of resistance for junctions which have been grouped by the breakdown voltage of the junction, distinguishing junctions as being above or below the mid-point of the bimodal distribution, 1.3 V. We show a difference in average resistance of the high vs. low breakdown junctions of $\sim$300 $\Omega$, a $\sim$4change. This further supports the idea that breakdown occurs at the thin-points in the barrier as anomalously thin points would result in a reduction in resistance and a reduction in breakdown voltage.  [^2] We of course cannot rule out a defect which both reduces the dielectric strength and increases the potential barrier height causing this discrepancy.
[…]
In Fig. 3 (e) we show the thickness distribution of two example junctions. We find that the distributions overlap quite well for counts at thinner distributions. Given that the junctions have similar resistances this is good evidence that the majority of the resistance is determined by the thinner portions of the barrier. We see this again in Fig. 3 (f) where we plot a cumulative conductance distribution as a function of junction area and find a large contribution to total conductance coming from a small fraction of the total area, irrespective of whether using a skewed lognormal distribution or a normally distributed thickness. Areas of locally elevated conductance are more prominent in the normally distributed junction, as is to be expected. As the thicker regions contribute little to the total conductance, our model has bad resolution on the thicker-side of the thickness distribution. The key feature of the log-normal probability distribution that our modeling does provide confidence for, is that the tails on the thinner-side of the distribution are suppressed more strongly than in the normal distribution. The observation of a small fraction of the junction contributing a large fraction of conductance agrees qualitatively with the conclusions from Refs. [39, 40]. Both studies arrive at this conclusion from considerations of barrier thickness measurements by (S)TEM, which we consider in the next section.
III Scanning Transmission Electron Microscopy
Figure 4: Experimental STEM-EDS imaging of an Al/AlOxAl barrier of the oxygen peak. (a-c) show heatmaps indicating the oxygen peak with the edges of the barrier as detected by Kernel integration overlaid to the maps. Different values of $\delta$ are used to define the kernel for the three maps (0, 0.2, 0.4 respectively). (d-f) show histograms of thicknesses as inferred from the different kernel integrations.
STEM includes a family of techniques which allow for sub-nanometre resolution imaging of materials with elemental contrast which are natural candidates to investigate questions about the thickness distribution of the barrier. These techniques can be applied to JJs where the barrier has been capped with a top electrode, i.e. after the full fabrication process has taken place. We first consider a protocol to measure variations in thickness in the projected 2D image. We do this considering images of the oxygen content in a barrier collected by STEM energy-dispersive X-ray spectroscopy (EDS) mapping. Imaging is performed on a lamella which is fabricated by focused ion beam cutting and lift out of a $\sim$30 nm thick cross section of the JJ. In Fig. 4 we present an example image of the oxygen composition. The edge of the barrier is identified by kernel integration using OpenCV [41]. We define two kernels to identify the two edges where each kernel has an asymmetry factor $\delta$. We describe the construction of the integration kernels in the supplemental information. Given that the oxygen content in the barrier increases over a finite length, the edge of the barrier could be defined at several points. By changing $\delta$ we systematically alter where the kernel integration protocol identifies the edge. In Fig. 4 (a-c) we show the same image with edges identified using 0, 0.2, 0.4 as values of $\delta$ respectively. In analysis we explored larger values of $\delta$ and found the edges were obviously within the junction, so care must be taken in selecting this asymmetry value.
We measure the thickness of the barrier by finding the distance between the two edges for each column of pixels. The average of the thickness from these different kernels matches previous measurements by TEM well giving an average thickness of $\sim$ 1.9 nm - 2.4 nm depending on the kernel [23]. We can also compute the distribution of thicknesses shown in the histograms Fig. 4 (d-f) and fit each of them both to normal and log-normal distributions. We routinely see that the log-normal distribution fits the thickness distribution better than the normal distribution. In the literature the results of this type of analysis have been interpreted as the statistical distribution of thicknesses inside a barrier, a claim we are explicitly not making here.
Arbitrary choices, such as the asymmetry factor in the kernels, can change what we infer about the barrier. We attribute this both to (i) details of the barrier including such as the oxygen content likely changing over a finite distance and the non-uniform barrier thickness and (ii) the STEM image being a projection which averages the barrier. As the barrier is grown on a bottom electrode which isn’t perfectly flat the imaging beam is at times intersecting both electrode and barrier. We wish to understand (i) from techniques such as STEM imaging as they will determine the behaviour of a JJ. However (ii) represents difficulties posed in trying to identify thickness variations in this material system by STEM. We suggest that we should report 1.9 nm $<\langle t\rangle<$ 2.4 nm, i.e. the variation in thickness implied by different kernels is what defines the error bars on the measurement of average barrier thickness.
In order to gain a deeper understanding of (ii) from above, we simulate STEM EDS images of oxygen content. We simulate the STEM image by building a 3D array of 0.1 nm3 voxels which represents the lamella used for STEM imaging. We use AFM maps to determine the topography of the barrier and then assign voxels in the barrier a value of 1, representing high oxygen content, voxels on the edge of the barrier a value of 0.5 and other voxels a value of 0, representing low oxygen content. Averaging values of voxels along a column normal to the lamella surface gives a value corresponding to the nominal oxygen content in that column and represents the projection formed by the electron beam experimentally. After projection we add Gaussian noise to the image representative of detector noise, we calibrate the centre and standard deviation of the Gaussian noise from a line trace of the experimental data in Fig. 4. We finally apply a Gaussian blur with radius 0.1 nm representative of the beam size in STEM experiments.
Figure 5: (a, b) two high resolution AFM scans of representative bottom leads (region indicated in Fig. 1 a) In (a, b) five 30 $\times$ 100 nm areas are marked by dashed lines. These regions are used to simulate STEM profiles shown in (c-g) assuming a 2 nm oxide barrier coating the AFM topography with the labels matching the AFM region to the STEM profile inserted into the boxes. (c-g) Edges of simulated oxygen content are detected using a symmetric kernels and shown overlaid, as well as measurements of average thickness and thickness standard deviations being noted in each figure. The labels (c - g) are in the same side of the scan region in the simulated STEM profile and the AFM regions.
We generate a series of barriers assuming that the thickness of the barrier is uniformly 2 nm across the junction. Using a uniform barrier thickness means that any thickness variation inferred from the simulated STEM image arises from (iii) and (iv) above, i.e. experimental nonidealities which could change what we infer from analysis of STEM images. We draw the bottom electrode topography from AFM measurements shown in Fig. 5 (a,b).  [^3] Noise at the $\sim$0.6 nm range is present in some regions of these AFM scans, which we found hard to exclude from measurements. This represents $\sim$10 % of the total topographic variation in these AFM maps and is not present across the whole scans and therefore we believe not a dominant contributor to the following analysis.. We infer the thickness of the barriers from the simulated STEM images (Fig. 5 c-g) using symmetric kernels ($\delta=0$) as described for Fig. 4 adjusting kernel sizes to account for the different pixel sizes. As expected, a perfectly uniform barrier appears to have varying thickness when viewed in projection due to roughness in the underlying barrier. We also find that the thickness values and thickness standard deviation change depending which AFM area is used to build the projection. All but one average thickness measurements are within 10 % of the barrier thickness input to the simulation, although the standard deviation of barrier thickness varies by a factor of 8$\sim$ 8$\times$ 8 ×. This suggests that STEM is a good technique to measure the average thickness of a barrier (reporting errors as described above) but calls into question the robustness of using STEM profiles to infer the thickness distribution in a junction. It also shows that multiple areas should be measured and inferring average thickness values from single STEM images may introduce errors. We therefore include additional STEM images with analogous processing to that performed in Fig. 4 in the supplemental material. These suggest similar thicknesses, but also include regions of topographic features which slightly increase the inferred average thickness.
Sometimes it is straightforward to identify regions where apparent thickness broadening occurs due to topographic variation in the bottom electrode (e.g. RHS of Fig. 5 c) where it’s possible to resolve two barriers superimposed on one another. Here the kernel edge detection gets confused shown by the discontinuities in the edge it identifies. In the supplemental material we show an example of experimental STEM imaging with these types of features, compared to a similar STEM simulation. In Fig. 5 (d) there is also a topographic gradient along the projection axis, but, in this instance, the topographic change is smaller in magnitude and so in the simulated STEM image it appears only as a broadening of the barrier. It is therefore not straightforwardly possible to systematically distinguish apparent broadening from topography variation and true barrier variation, although it is possible to exclude areas where the topographic broadening effects are largest.
Comparing experimental (Fig. 4) and simulated (Fig. 5) STEM images, barriers appear rougher in the experimental STEM data than the simulated STEM. The topography of the electrode used to simulate STEM images may be artificially smoothed due to the AFM tip radius of 2 nm. This would mean that features of the size of apparent roughness seen in Fig. 4 would not be apparent with this AFM tip. The barriers we measure here are also oxidised which can change the topography of the surface. While there may be features sufficiently small that they are not resolved by AFM, both AFM maps in Fig. 5 have topographic features are large relative to the AFM resolution, and as we show, couple importantly into the inferred STEM projections. Using STM may allow higher resolution topographic maps to perform more robust inferences of STEM profiles, and can also be performed in situ before barrier oxidation modifies the topography. These more detailed maps might find uses, for instance in reverse Monte-Carlo approaches [43] to extract quantitative information about the barrier thickness and thickness distributions from STEM.
IV Conclusions
Technique
Simmons
Model Fit
Monte-Carlo
Simulations
STEM-EDS
Variable -
Potential Barrier Height (eV)
Only log-normal distribution
Kernel Asymmetry ($\delta$)
0.8 1.0 1.22 0 0.2 0.4
Thickness
0.73 $\pm$ 0.02 2.10 $\pm$ 0.55 1.45$\pm$ 0.38 1.02 $\pm$ 0.21 2.39 $\pm$ 0.32 2.13 $\pm$ 0.22 1.94 $\pm$ 0.17
Table 1: A summary of different barrier thicknesses extracted throughout this manuscript. Fitting the Simmons model returns a single-valued thickness and the error represents the standard deviation from fitting 598 IVs. Monte-Carlo simulations give thickness by fixing the area of a junction and allowing the barrier thickness to vary. The thickness value is the average thickness whereas the errors represent a genuine thickness distribution in the barrier. STEM-EDS measurements are direct measurements of thickness by atomic scale microscopy. As shown here, the thickness of the barrier is sensitive to image processing (i.e. choice of $\delta$). Errors here represent standard deviations of inferred average thickness for a specific choice of $\delta$ but do not represent thickness distributions. Without an un-biased approach to image processing, the true error on thickness measurements is likely the variation in thickness between different values of $\delta$.
In this article we have aimed to understand oxide barriers in JJs due to their technological importance in superconducting qubits. A summary of different thicknesses inferred within this paper is shown in Table 1. We fit the Simmons model to non-linear IVs to extract a single-valued thickness value to each junction $\sim$0.75 nm. We build Monte-Carlo protocols to simulate the IV properties of barriers with multi-valued thicknesses drawn from example statistical distributions Here we find that the distribution of thicknesses present in the junction, results in IVs that, when fitted, would imply a single-value thickness smaller than the average thickness. This is because conductivity is exponentially dependent on the barrier thickness, so any thinner-than-average regions contribute more than similarly thicker-than-average regions. These simulations indicate a barrier that is $\sim$1 nm thick with a standard deviation of $\sim$0.2 nm thick, however using different barrier heights, we infer different thickness distributions. The rectangular potential barrier used in the Simmons model likely results in an under-estimate of thickness relative to more realistic, smoothly increasing barriers. By considering the Monte Carlo simulations relative to STEM measurements we constrain the barrier height to the range 0.8 eV ϕless-than-or-similar-toabsentitalic-$\lesssim\phi\lesssim$ 1.22 eV.
Using breakdown measurements we can probe the tails of thickness distributions assuming that breakdown occurs at the thinnest point or points of the barrier. Considering normal and skewed log-normal distributions of barrier thickness, we show that the breakdown statistics are better described by a skewed distribution. We show that the tails of the thickness distribution contribute substantially to the resistance of the device with a $\sim$4 % difference in resistance between the 93 % lowest- and 7 % highest-breakdown voltage junctions. We also show that normally distributed barrier thicknesses give rise to larger barrier-induced spread than skewed distributions.
We then consider STEM measurements probing the oxygen content of barriers as a technique to directly measure the thickness of barriers. We show measurements of the barrier thickness from 1.9 to 2.4 nm depending on how the edge of the barrier is determined with a standard deviation not dissimilar to those inferred from the Monte Carlo simulations of the barrier. We suggest an approach to define errors on barrier thickness measurements, performing kernel integration with different kernels to extract different barrier edges and comparing the variation in the average thickness reported. These error bars should not be misinterpreted as the standard deviation of the barrier thickness.
Comparing experimental measurements to simulations of STEM data, we show that inferring average thicknesses from STEM images is reasonable, subject to errors associated with identifying the edge of the barrier. However, inferring variations in the thickness of the barrier from STEM images is not a robust process. The roughness commonly seen on bottom electrodes, when projected into 2D, also creates apparent thickness variations. We see wider variations in barrier thickness measurements experimentally than when simulating a perfectly uniform barrier, lending support to the idea that the barrier thickness does indeed vary across the junction but inferring the statistics of this variation from STEM data alone remains a challenge. Most evidence in the literature about thickness distributions of barriers and that this is normally distributed arises from this type of processing of STEM or TEM data for instance [23]. While these works were substantial increments to the understanding of the oxide barriers in JJs, continuing to use these techniques and inferences should be avoided, for instance in computations of Josephson harmonics based on normal thickness distributions [18]. We show here that this type of inference over-states the certainty of thickness distributions, and indeed our most sensitive probe of the tails of our distribution imply some form of skewed thickness distribution contradicting the conclusions drawn from STEM.
Acknowledgements.
The authors thank the OQC Fabrication team for discussions and contributions toward our Josephson junction manufacturing process. The STEM imaging was taken in LEMAS center in the University of Leeds, thanks to John Harrington and Zabeada Aslam for FIB sectioning and STEM operation. We thank the Royal Holloway University of London SuperFab Facility for their support.